# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ I.4 Spatial propagation of action potentials in coupled FitzHugh-Nagumo models
As shown in Fig. 2B, the FHN equation shows both excitable and oscillatory behavior, thus capturing typical neuronal dynamics. In reality, such action potentials propagate along the axon. Therefore, diffusion in space was added to both the HH model and the FHN model:
$\begin{array}[]{ @{} l @{} }u_{t}=D_{u}\Delta u-u^{3}+u-v,\\[3.0pt] v_{t}=D_{v}\Delta v+\varepsilon(u-bv+a).\end{array}$ (10)
@@
$\begin{array}[]{ @{} l @{} }{u_{i}}_{t}=-u_{i}^{3}+u_{i}-v_{i}+\sum_{j}^{n}p(u_{j},v_{j})\\[3.0pt] {v_{i}}_{t}=\varepsilon(u_{i}-bv_{i}+a)+\sum_{j}^{n}q(u_{j},v_{j})\end{array}.$ (11)
@@ I.5 Beyond neuroscience: diverse applications of the FitzHugh-Nagumo model
Despite its simplicity, the FHN model captures many of the essential features of neuronal dynamics, and has been used to study a wide range of phenomena in neuroscience FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Longtin (1993); Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Nagumo et al. (1962); Bressloff (2014); Murray (2001), including the synchronization of neural activity Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Al Beattie et al. (2024), the dynamics of neural networks FitzHugh (1961); Izhikevich (2007a); Upadhyay and Iyengar (2021), and the emergence of complex spatiotemporal patterns in neural systems Murray (2001) (Fig. 1 D). However, thanks to its generic dynamical behavior, the FHN model finds applications in various fields beyond neuroscience. In Fig. 3, we provide an overview of the main fields where the FHN model has helped to advance our understanding of those systems. We highlight electronics, optics, and biological systems. Given its significant role in cardiac modeling, the FHN model’s application to cardiac systems is discussed separately. Motivated by its broad applicability and generic dynamics, the FHN system has triggered a large body of theoretical work, characterizing bifurcations and its various dynamical behaviors. Fig. 3 illustrates the work in the following domains:
Neuroscience: The FHN model, initially developed in neuroscience, has played a pivotal role in understanding the complex dynamics of neurons and neural networks FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Longtin (1993); Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Nagumo et al. (1962); Bressloff (2014); Murray (2001); Al Beattie et al. (2024). By distilling complex neuronal behaviors into simple equations, the FHN model has made it easier to analyze and simulate neural dynamics. It offers a comprehensive view of solutions, providing geometric insights into crucial biological phenomena like neuronal excitability and spike generation mechanisms.
@@
Moreover, the FHN model has proven invaluable in studying neuronal interactions within networks. By linking multiple FHN units, researchers have gained deeper insights into network communication, synchronization, and the generation of coordinated activity patterns. This has greatly enhanced our understanding of phenomena like neuronal oscillations and information processing in the brain. Furthermore, the FHN model has laid the groundwork for the development of more sophisticated neuronal models, serving as a platform for incorporating additional biological intricacies. By leveraging insights from the FitzHugh-Nagumo framework, researchers continue to refine our understanding of neuronal function and dysfunction in both health and disease.
@@
Among the various models used to study cardiac cells, the FHN model stands out as one of the simplest yet most widely studied for capturing the general dynamical features of cardiac cells Keener (1988); Clayton and Panfilov (2008); Keener and Sneyd (2009); Van Capelle and Durrer (1980); Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000); Alonso et al. (2016); Bär and Eiswirth (1993b); Alonso et al. (2003). In addition to the FHN model, several other generic models are used to reproduce the qualitative properties of cardiac excitation. The Barkley model, which is very similar to the FHN model, is widely used in numerical studies due to its efficient numerical implementation Dowle et al. (1997); Barkley (1991). The Bär–Eiswirth model extends the Barkley model and was introduced specifically to study spiral breakup Bär and Eiswirth (1993b). Similarly, the Puschino model is extensively utilized in numerical simulations of excitable media Panfilov and Rudenko (1987). The Karma model, another variant, was developed to demonstrate spiral wave breakup due to alternans Karma (1993). It generates an action potential with a more realistic fast upstroke and slow recovery compared to the standard FHN model. Additionally, the Aliev-Panfilov model adapts the FHN framework to capture the rate-dependent variations in action potential duration Aliev and Panfilov (1996).
In cardiac systems, many models are employed to investigate the existence and dynamics of spirals, which are wave phenomena arising from anisotropy or spatial defects and characterized by a core and rotation frequency Keener (1986b). Electrical impulses, denoted by the variable $u$, propagate in a wave-like manner in the heart, and it has been observed that ventricular fibrillations result from the collision of two spirals rotating in opposite directions Gray et al. (1995); Jalife (1990). The FHN model is frequently utilized to simulate such behavior and analyze the effects of spiral breakup Davidenko et al. (1995); Winfree (1991a); Clayton and Panfilov (2008); Keener and Sneyd (2009); Panfilov and Hogeweg (1993); Franzone et al. (2006); Aliev and Panfilov (1996); Van Capelle and Durrer (1980).
@@
In the context of cellular processes, the FHN model has been employed to describe the oscillatory dynamics of the activity of the cyclin B–cyclin-dependent kinase 1 (Cdk1) complex during the cell division cycle Parra-Rivas et al. (2023a); Gelens et al. (2014); Nolet et al. (2020a); Rombouts and Gelens (2020); Gelens et al. (2015); Tsai et al. (2008). Additionally, it has been instrumental in studying oscillations in the concentration of free cytosolic Ca2+, a crucial cellular control mechanism Duffy et al. (1998). Beyond cellular biology, the FHN model has found applications in diverse organisms. In bacteria, it has been utilized to investigate polarization dynamics in response to electrical stimuli, akin to neuronal responses Stratford et al. (2019). In plants, FHN-like models have served as cell signaling models for pulse-like jasmonate responses Chiangga et al. (2016).
@@
The coupling of FHN units in networks has shed light on various physiological phenomena. For instance, in mammalian pancreas $\beta$ cells, coupled FHN units elucidate how global oscillations emerge despite the absence of a pacemaker region Cartwright (2000); Scialla et al. (2021). Similarly, in plant systems, coupling FHN models with photosynthesis models aids in understanding the impact of environmental variables such as temperature Hartzell et al. (2015). Spatially extended systems described by diffusively coupled FHN equations have provided insights into wave propagation during mitosis in Xenopus frog embryos Gelens et al. (2014); Nolet et al. (2020a); Rombouts and Gelens (2020); Gelens et al. (2015), as well as calcium waves triggered by fertilization Sneyd et al. (2017b). The FHN model has also been applied to study propagating action potentials in vascular plant tissues Vodeneev et al. (2011); Sukhov et al. (2011) and morphoelastic waves driving the locomotion of soft robots Miller and Dunkel (2020). In ecology, adding cross-diffusive effects to the FHN model has enabled studies on pattern formation in living systems, such as chemotactic movement in Escherichia coli and interactions in predator-prey dynamics Gambino et al. (2022, 2023); Zemskov et al. (2011); Lee (1997).
@@
Electronic and optical systems: The significance of understanding the dynamics of neurons and neural networks has long been acknowledged, both for unraveling the complexities of the human brain and for applying analogous concepts to machine learning applications. Consequently, researchers have endeavored to emulate neuron cells using electronic circuits. Many of these circuits have been specifically devised to replicate the dynamics of the FHN model Linares-Barranco et al. (1991); Keener (1983); Khakipoor et al. (2022); Xu et al. (2020); Bao et al. (2019); Zhang and Liao (2017); Liu et al. (2020); Li et al. (2021); Binczak et al. (2006, 2003); Takembo et al. (2019).
@@
In recent years, advancements in integrating optoelectronic components onto photonic integration platforms have spurred extensive investigations into ultrafast artificial neural networks for information processing Shastri et al. (2021). All-optical spiking neurons have been successfully demonstrated using semiconductor lasers Gelens et al. (2010); Coomans et al. (2011); Selmi et al. (2014); Nahmias et al. (2013). Furthermore, neuronal dynamics described by the FHN system have been realized through (electro-)optical implementations García-Ojalvo and Schimansky-Geier (1999); Rosin et al. (2011); Romeira et al. (2016).
@@
Mathematics and physics: While the FHN model was specifically developed to describe neuronal dynamics, it has served more generally as a cornerstone in understanding the rich dynamical behaviors exhibited by excitable systems through dynamical systems analysis Rocsoreanu et al. (2012); Wang et al. (2009); Vaidyanathan (2015); Hoff et al. (2014); Burić and Todorović (2003); Lei and Yang (2021); Brandibur and Kaslik (2022); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Campbell and Waite (2001); Toral et al. (2003a); Ginoux and Llibre (2015); Toral et al. (2003b); Yu et al. (2021); Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Rotstein et al. (2012); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Lindner (2004); Wiesenfeld et al. (1994); Desroches et al. (2012, 2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005b); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007); Schöll et al. (2009); Perc (2005a); Steur et al. (2009); Makarov et al. (2001); Zaks et al. (2005); Gambino et al. (2022); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Pankratova et al. (2005); Hastings (1982); Perlikowski et al. (2010); Baer and Erneux (1992, 1986); Schmidt et al. (2017); Lin et al. (2004); García-Ojalvo and Schimansky-Geier (1999); Saha and Feudel (2017); Bode (1997); Gerster et al. (2020); Tobias and Knobloch (1998); Yanagita et al. (2005); Shoji et al. (2007); Kostova et al. (2004). By representing the essential characteristics of excitable cells, the FHN model provides a simplified yet powerful framework for exploring a wide range of dynamical phenomena. Through bifurcation analysis, researchers have uncovered complex patterns of behavior emerging from the model’s equations. These analyses reveal how changes in parameters, such as excitability thresholds or coupling strengths, can lead to transitions between different dynamical regimes, including fixed points, limit cycles, and chaotic behavior. Moreover, stability analysis techniques enable the identification of critical points where small perturbations can lead to qualitatively different system behaviors, shedding light on the robustness and sensitivity of excitable systems.
@@
Furthermore, phase-plane analysis has been instrumental in elucidating the dynamics of the FHN model by visualizing trajectories in its phase space. By plotting the evolution of variables such as membrane potential and recovery variable, researchers can gain insights into the underlying mechanisms driving excitability, oscillations, and wave propagation. Through phase-plane analysis, the FHN model’s behaviors, such as action potential generation and refractoriness, can be understood in terms of the system’s underlying dynamics. Additionally, techniques such as nullclines analysis provide a geometric understanding of the model’s behavior, revealing regions of parameter space associated with different dynamical regimes.
@@
Other: Additional examples not mentioned previously encompass chemical reaction dynamics Kuramoto et al. (1998); Lin et al. (2004); Tobias and Knobloch (1998), the dynamics of elastic excitable media Cartwright et al. (1999), or the study of computational algorithms Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994). In exploring chemical reaction dynamics, the FHN model has been instrumental in investigating the emergence of chemical turbulence through reaction and diffusion processes Kuramoto et al. (1998); Tobias and Knobloch (1998). One mechanism for generating chemical turbulence involves spiral waves that undergo breakup, a phenomenon initially demonstrated in the context of the Belousov-Zhabotinsky chemical reaction-diffusion system Ouyang and Flesselles (1996), and subsequently studied extensively within the framework of the FHN model Tobias and Knobloch (1998). Furthermore, the FHN system has been employed to comprehend spatial patterns in a forced FitzHugh-Nagumo reaction-diffusion model, reproducing dynamics akin to those observed in the Belousov-Zhabotinsky system Lin et al. (2004).
@@
In the context of elastic excitable media, the FHN-equivalent Burridge–Knopoff model has been applied to characterize the frictional sliding dynamics in earthquake fault systems influenced by viscous friction Cartwright et al. (1999). Due to the limited number of found references in these fields, they were not included as a separate category in the table depicted in Fig. 3.

### v4

@@ I.4 Spatial propagation of action potentials in coupled FitzHugh-Nagumo models
While the excitable and oscillatory behavior in the FHN equation capture typical neuronal dynamics (Fig. 2B), in reality, the action potentials also propagate along the axon. Therefore, diffusion in space was added to both the HH model and the FHN model:
$\begin{array}[]{ @{} l @{} }u_{t}=D_{u}\Delta u-u^{3}+u-v,\vspace{3pt}\\ v_{t}=D_{v}\Delta v+\varepsilon(u-bv+a).\end{array}$ (10)
@@
$\begin{array}[]{ @{} l @{} }{u_{i}}_{t}=-u_{i}^{3}+u_{i}-v_{i}+\sum_{j}^{n}p(u_{j},v_{j}),\vspace{3pt}\\ {v_{i}}_{t}=\varepsilon(u_{i}-bv_{i}+a)+\sum_{j}^{n}q(u_{j},v_{j}).\end{array}$ (11)
@@
Figure 3: Overview of FitzHugh-Nagumo model applications This table organizes selected literature related to different formulations of the FitzHugh-Nagumo equations, as outlined in Eq. 8, Eq. 10, and Eq. 11. Each listed study is categorized by its application domain and primary focus. It is important to note that, alongside the conventional FitzHugh-Nagumo models, certain studies incorporate variations like piecewise linear models or three-dimensional adaptations to meet particular investigative requirements.
@@ I.5 Beyond neuroscience: diverse applications of the FitzHugh-Nagumo model
Despite its simplicity, the FHN model captures many of the essential features of neuronal dynamics, and has been used to study a wide range of phenomena in neuroscience FitzHugh (1961); Izhikevich (2006); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Longtin (1993); Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Nagumo et al. (1962); Bressloff (2014); Murray (2001), including the synchronization of neural activity Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Al Beattie et al. (2024), the dynamics of neural networks FitzHugh (1961); Izhikevich (2006); Upadhyay and Iyengar (2021), and the emergence of complex spatio-temporal patterns in neural systems Murray (2001) (Fig. 1D). However, thanks to its generic dynamical behavior, the FHN model finds applications in various fields beyond neuroscience. In Fig. 3, we provide an overview of the main fields where the FHN model has helped to advance our understanding of those systems. We highlight electronics, optics, and biological systems. Given its significant role in cardiac modeling, the FHN model’s application to cardiac systems is discussed separately. Motivated by its broad applicability and generic dynamics, the FHN system has triggered a large body of theoretical work, characterizing bifurcations and its various dynamical behaviors. Fig. 3 illustrates the work in the following domains:
Neuroscience: The FHN model, initially developed in neuroscience, has played a pivotal role in understanding the complex dynamics of neurons and neural networks FitzHugh (1961); Izhikevich (2006); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Longtin (1993); Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Nagumo et al. (1962); Bressloff (2014); Murray (2001); Al Beattie et al. (2024). By describing complex neuronal behaviors with simple equations, the FHN model has made it easier to analyze and simulate neural dynamics. It offers a comprehensive view of solutions, providing geometric insights into crucial biological phenomena like neuronal excitability and spike generation mechanisms.
@@
Moreover, the FHN model has proven invaluable in studying neuronal interactions within networks. By linking multiple FHN units, researchers have gained deeper insights into network communication, synchronization, and the generation of coordinated activity patterns. This has greatly enhanced our understanding of phenomena like neuronal oscillations and information processing in the brain. Furthermore, the FHN model has laid the groundwork for the development of more sophisticated neuronal models, serving as a platform for incorporating additional biological complexity. By leveraging insights from the FitzHugh-Nagumo framework, researchers continue to refine our understanding of neuronal function and dysfunction in both health and disease.
@@
Among the various models used to study cardiac cells, the FHN model stands out as one of the simplest yet most widely studied for capturing the general dynamical features of cardiac cells Keener (1988); Clayton and Panfilov (2008); Keener and Sneyd (2009); Van Capelle and Durrer (1980); Franzone et al. (2006); Davidenko et al. (1995); Courtemanche et al. (1990); Keener and Sneyd (2009); Aliev and Panfilov (1996); Biktashev (2002); Pertsov et al. (1993); Nash and Panfilov (2004); Watanabe et al. (2004); Sermesant et al. (2006); Cherubini et al. (2008); Glass and Josephson (1995); Le and Sotiropoulos (2013); Panfilov and Keener (1995); Nagai et al. (2000); Alonso et al. (2016); Bär and Eiswirth (1993a); Alonso et al. (2003); Ben Abid et al. (2024). In addition to the FHN model, several other generic models are used to reproduce the qualitative properties of cardiac excitation. The Barkley model, which is very similar to the FHN model, is widely used in numerical studies due to its efficient numerical implementation Dowle et al. (1997); Barkley (1991). The Bär-Eiswirth model extends the Barkley model and was introduced specifically to study spiral breakup Bär and Eiswirth (1993a). Similarly, the Puschino model is extensively utilized in numerical simulations of excitable media Panfilov and Rudenko (1987). The Karma model, another variant, was developed to demonstrate spiral wave breakup due to alternans Karma (1993). It generates an action potential with a more realistic fast upstroke and slow recovery compared to the standard FHN model. Additionally, the Aliev-Panfilov model adapts the FHN framework to capture the rate-dependent variations in action potential duration Aliev and Panfilov (1996).
In cardiac systems, many models are used to investigate the existence and dynamics of spirals, which are wave phenomena arising from anisotropy or spatial defects and characterized by a core and rotation frequency Keener (1986a). Electrical impulses, denoted by the variable $u$, propagate in a wave-like manner in the heart, and it has been observed that ventricular fibrillations result from the collision of two spirals rotating in opposite directions Gray et al. (1995); Jalife (1990). The FHN model is frequently utilized to simulate such behavior and analyze the effects of spiral breakup Davidenko et al. (1995); Winfree (1991a); Clayton and Panfilov (2008); Keener and Sneyd (2009); Panfilov and Hogeweg (1993); Franzone et al. (2006); Aliev and Panfilov (1996); Van Capelle and Durrer (1980).
@@
In the context of cellular processes, the FHN model has been employed to describe the oscillatory dynamics of the activity of the cyclin B-cyclin-dependent kinase 1 (Cdk1) complex during the cell division cycle Parra-Rivas et al. (2023a); Gelens et al. (2014); Nolet et al. (2020a); Rombouts and Gelens (2020); Gelens et al. (2015); Tsai et al. (2008). Additionally, it has been instrumental in studying oscillations in the concentration of free cytosolic Ca2+, a crucial cellular control mechanism Duffy et al. (1998). Beyond cellular biology, the FHN model has found applications in diverse organisms. In bacteria, it has been used to investigate polarization dynamics in response to electrical stimuli, akin to neuronal responses Stratford et al. (2019). In plants, FHN-like models have served as cell signaling models for pulse-like jasmonate responses Chiangga et al. (2016).
@@
The coupling of FHN units in networks has shed light on various physiological phenomena. For instance, in mammalian pancreas $\beta$ cells, coupled FHN units elucidate how global oscillations emerge despite the absence of a pacemaker region Cartwright (2000); Scialla et al. (2021). Similarly, in plant systems, coupling FHN models with photosynthesis models aids in understanding the impact of environmental variables such as temperature Hartzell et al. (2015). Spatially extended systems described by diffusively coupled FHN equations have provided insights into wave propagation during mitosis in Xenopus frog embryos Gelens et al. (2014); Nolet et al. (2020a); Rombouts and Gelens (2020); Gelens et al. (2015), as well as calcium waves triggered by fertilization Sneyd et al. (2017a). The FHN model has also been applied to study propagating action potentials in vascular plant tissues Vodeneev et al. (2011); Sukhov et al. (2011) and morphoelastic waves driving the locomotion of soft robots Miller and Dunkel (2020). In ecology, adding cross-diffusive effects to the FHN model has enabled studies on pattern formation in living systems, such as chemotactic movement in Escherichia coli and interactions in predator-prey dynamics Gambino et al. (2022, 2023); Zemskov et al. (2011); Lee (1997).
@@
Electronic and optical systems: The significance of understanding the dynamics of neurons and neural networks has long been acknowledged, both for unraveling the complexities of the human brain and for applying analogous concepts to machine learning applications. Consequently, researchers have endeavored to emulate neuron cells using electronic circuits. Many of these circuits have been specifically devised to replicate the dynamics of the FHN model Linares-Barranco et al. (1991); Keener (1983); Khakipoor et al. (2022); Xu et al. (2020); Bao et al. (2019); Zhang and Liao (2017); Liu et al. (2020); Barland et al. (2003); Li et al. (2021); Binczak et al. (2006, 2003); Takembo et al. (2019).
@@
In recent years, advancements in integrating optoelectronic components onto photonic integration platforms have spurred extensive investigations into ultrafast artificial neural networks for information processing Shastri et al. (2021). All-optical spiking neurons have been successfully demonstrated using semiconductor lasers Gelens et al. (2010); Coomans et al. (2011); Selmi et al. (2014); Nahmias et al. (2013). Furthermore, neuronal dynamics described by the FHN system have been realized through (electro-)optical implementations García-Ojalvo and Schimansky-Geier (1999); Rosin et al. (2011); Romeira et al. (2016); Huang et al. (2024a).
@@
Mathematics and physics: While the FHN model was specifically developed to describe neuronal dynamics, it has served more generally as a cornerstone in understanding the rich dynamical behaviors exhibited by excitable systems through dynamical systems analysis Rocsoreanu et al. (2012); Wang et al. (2009); Vaidyanathan (2015); Hoff et al. (2014); Burić and Todorović (2003); Lei and Yang (2021); Brandibur and Kaslik (2022); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Campbell and Waite (2001); Toral et al. (2003a); Ginoux and Llibre (2015); Toral et al. (2003b); Yu et al. (2021); Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Ohta et al. (1989); Iqbal et al. (2017); Elphick et al. (1990); Meron (1992); Petrich and Goldstein (1994); Hagberg and Meron (1994a); Goldstein et al. (1996); Vasiev (2004); Nekhamkina and Sheintuch (2006); Rotstein et al. (2012); Zheng and Shen (2015); Tsyganov et al. (2014); Rombouts and Gelens (2020); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Zaritski et al. (2004); Kuznetsov et al. (2017); Shepelev et al. (2017); Browne et al. (2008); Lindner et al. (2004); Wiesenfeld et al. (1994); Desroches et al. (2012, 2008); Guckenheimer and Kuehn (2010, 2009); Perc (2005a); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Semenova et al. (2016); Schülen et al. (2022); Schöll (2016b); Omelchenko et al. (2015); Perc (2007); Schöll et al. (2009); Perc (2005b); Steur et al. (2009); Makarov et al. (2001); Zaks et al. (2005); Gambino et al. (2022); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Métens et al. (1997); Zemskov and Epstein (2010); Pankratova et al. (2005); Hastings (1982); Perlikowski et al. (2010); Baer and Erneux (1992, 1986); Schmidt et al. (2017); Lin et al. (2004); García-Ojalvo and Schimansky-Geier (1999); Saha and Feudel (2017); Bode (1997); Gerster et al. (2020); Tobias and Knobloch (1998); Yanagita et al. (2005); Shoji et al. (2007); Kostova et al. (2004). By representing the essential characteristics of excitable cells, the FHN model provides a simplified yet powerful framework for exploring a wide range of dynamical phenomena. Through bifurcation analysis, researchers have uncovered complex patterns of behavior emerging from the model’s equations. These analyses reveal how changes in parameters, such as excitability thresholds or coupling strengths, can lead to transitions between different dynamical regimes, including fixed points, limit cycles, and chaotic behavior. Moreover, stability analysis techniques enable the identification of critical points where small perturbations can lead to qualitatively different system behaviors, shedding light on the robustness and sensitivity of excitable systems.
@@
Furthermore, phase-plane analysis has been instrumental in elucidating the dynamics of the FHN model by visualizing trajectories in its phase space. By plotting the evolution of variables such as membrane potential and recovery variable, one can gain insights into the underlying mechanisms driving excitability, oscillations, and wave propagation. Through phase-plane analysis, the FHN model’s behaviors, such as action potential generation and refractoriness, can be understood in terms of the system’s underlying dynamics. Additionally, techniques such as nullclines analysis provide a geometric understanding of the model’s behavior, revealing regions of parameter space associated with different dynamical regimes.
@@
Other: Additional examples not mentioned previously encompass chemical reaction dynamics Kuramoto et al. (1998); Lin et al. (2004); Tobias and Knobloch (1998), the dynamics of elastic excitable media Cartwright et al. (1999), or the study of computational algorithms Moayeri et al. (2020); Dehghan and Narimani (2018); Fuselier and Wright (2013); Rogers and McCulloch (1994). In exploring chemical reaction dynamics, the FHN model has been instrumental in investigating the emergence of chemical turbulence through reaction and diffusion processes Kuramoto et al. (1998); Tobias and Knobloch (1998). One mechanism for generating chemical turbulence involves spiral waves that undergo breakup, a phenomenon initially demonstrated in the context of the Belousov-Zhabotinsky chemical reaction-diffusion system Ouyang and Flesselles (1996), and subsequently studied extensively within the framework of the FHN model Tobias and Knobloch (1998). Furthermore, the FHN system has been employed to understand spatial patterns in a forced FitzHugh-Nagumo reaction-diffusion model, reproducing dynamics akin to those observed in the Belousov-Zhabotinsky system Lin et al. (2004).
@@
In the context of elastic excitable media, the FHN-equivalent Burridge-Knopoff model has been applied to characterize the frictional sliding dynamics in earthquake fault systems influenced by viscous friction Cartwright et al. (1999). Due to the limited number of found references in these fields, they were not included as a separate category in the table depicted in Fig. 3.